# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

### v2

@@ Abstract
UAV-View Geo-Localization (UVGL) aims to achieve accurate localization of unmanned aerial vehicles (UAVs) by retrieving the most relevant GPS-tagged satellite images. However, existing methods heavily rely on pre-paired UAV-satellite images for supervised learning. Such dependency not only incurs high annotation costs but also severely limits scalability and practical deployment in open-world UVGL scenarios. To address these limitations, we propose an end-to-end self-supervised UVGL method. Our method leverages a shallow backbone network to extract initial features, employs clustering to generate pseudo labels, and adopts a dual-path contrastive learning architecture to learn discriminative intra-view representations. Furthermore, our method incorporates two core modules, the dynamic hierarchical memory learning module and the information consistency evolution learning module. The dynamic hierarchical memory learning module combines short-term and long-term memory to enhance intra-view feature consistency and discriminability. Meanwhile, the information consistency evolution learning module leverages a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, thereby improving cross-view feature alignment. To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced, which refines the quality of pseudo supervision. Our method ultimately constructs a unified cross-view feature representation space under self-supervised settings. Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods. These results validate the effectiveness and generalizability of our method, highlighting its potential for real-world deployment in UVGL. Our code is available at https://github.com/ISChenawei/DMNIL.
@@
UAV-View Geo-Localization, Self-Supervised Learning, Contrastive Learning.I Introduction
@@
UAV-View Geo-Localization (UVGL) aims to achieve high-precision localization of UAV in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to UAV-captured near-ground view images [1]. This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response. In recent years, existing UVGL methods have primarily relied on supervised learning using strictly paired UAV-satellite images [2, 3, 4]. Although these supervised methods have achieved encouraging success, they also reveal several inherent limitations: (1) acquiring instance-level paired UAV-satellite images is highly costly [5], leading to increased training overhead; (2) large amounts of unpaired cross-view image resources remain underutilized; and (3) when applied to new scenarios, these methods often require recollection and re-pairing UAV-satellite images for training to adapt to domain shifts. These limitations significantly hinder the scalability and practical deployment of UVGL in real-world scenarios. Thereby, it is necessary and of great significance to explore a self-supervised UVGL method that can effectively mine latent cross-view correlations from large-scale unpaired UAV-satellite images.
In recent years, self-supervised person re-identification (ReID) methods [6, 7] have achieved remarkable progress, offering crucial theoretical foundations for UVGL. Typically, These methods employ clustering algorithms [8, 9] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning. However, the quality of these pseudo-labels is heavily dependent on the discriminative capability of backbone network. As illustrated in Fig. 1, compared with conventional ReID, UVGL faces additional challenges. These factors make it more difficult for the backbone to extract discriminative features, as illustrated in Fig. 2. When pseudo-labels contain a high error rate and the model lacks the capacity to capture intra-view and cross-view consistency and discriminative features, the self-supervised learning process becomes prone to catastrophic negative optimization.
@@
In this paper, we propose a novel end-to-end self-supervised UVGL method. A lightweight, shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels. Combined with a dual-path contrastive learning strategy [17, 18], these components form the baseline network of our self-supervised method for UVGL to learn intra-view feature representations. Building upon this baseline, we proposed the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method. The DHML module captures the relationship between short-term and long-term memory to learn the discriminability of intra-view feature representations. The ICEL module constructs a neighborhood-driven dynamic constraint mechanism based on feature distributions. By integrating neighborhood consistency constraints with mutual information optimization, it effectively aligns cross-view feature representations within a shared embedding space, promoting both consistency and discriminative learning. Additionally, to improve the reliability of pseudo-labels, we introduce a pseudo-label enhancement (PLE) strategy, which contributes to a more stable and effective self-supervised feature learning process. The main contributions can be summarized as follows:
@@
We propose an end-to-end self-supervised method for UVGL. Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to effectively capture and enhance feature consistency and discriminability for both intra-view and cross-view representation learning.
@@
We propose a dynamic memory hierarchical learning module that captures the relationships among different levels of memory to adaptively model intra-class and inter-class feature dynamics within the intra-view, thereby enhancing the discriminability of feature representations.
@@
we propose a novel information consistency evolution learning module which combines neighborhood consistency and mutual information optimization to learn consistent and discriminative cross-view representations.
@@
Extensive experiments demonstrate that our method, utilizing only a lightweight backbone, achieves state-of-the-art performance on three UVGL benchmarks, outperforming even certain supervised methods.
@@
The remainder of this paper is organized as follows. Section II systematically reviews prior research. In Section III, the proposed self-supervised UVGL method is presented in detail. The experimental results are reported and analyzed in Section IV. Finally, conclusions are outlined in Section V.
@@ II RELATED WORKS
This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised UVGL method. Although this work focuses on UVGL as a sub-task of cross-view geo-localization (CVGL), we first review the overall progress in supervised CVGL. We then discuss recent advances in self-supervised CVGL methods. Finally, we summarize main methodes in self-supervised ReID.
@@ II-A Supervised Cross-View Geo-localization
CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images. Early research in CVGL primarily focused on the ground-view CVGL [19, 20, 21], leading to the development of several benchmark datasets, such as CVUSA [22], CVACT [23], and VIGOR [24]. To address the challenge of ground-view CVGL, researchers have proposed various methods. CVM-Net [25] employed NetVLAD [26] to encode images as global descriptors, thereby reducing visual discrepancies caused by viewpoint variations. TransGeo [27] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization. GeoDTR [20] leveraged a Transformer-based [28] feature extractor to disentangle geometric features, effectively addressing perspective shifts. Sample4Geo [29] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
With the continuous development and application of drone technology, UVGL has gradually become an important research direction and achieves significant progress [30, 31, 32]. Several benchmark datasets for UVGL have been proposed, such as University-1652[12], SUES-200, [33] and DenseUAV [34], which have driven further research in this field. In this scenario, a variety of methods have been proposed by researchers. LPN [35] mined fine-grained features through local pattern partitioning, improving the ability to capture local details. IFS [36] utilized a multi-branch strategy to effectively fuse global and local features, enhancing the diversity and robustness of feature representations. CAMP [30] improved cross-view matching accuracy by contrastive attribute mining and position-aware partitioning, while DAC [4] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints. MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve feature representation and alignment.
Nevertheless, the remarkable performances exhibited by these UVGL methods require extensive pre-processed UAV-satellite image pairs. In this work, we pivot our focus to the realm of self-supervised UVGL, where prior pairing relationships are absent, presenting important applications for real-world UVGL deployments.
@@ II-B Self-Supervised Cross-View Geo-localization
Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for real-world, researchers have begun to explore self-supervised CVGL methods. The work in [37] proposed a self-supervised method that leveraged correspondence-free projection to transform ground panorama images into bird’s-eye view images. It also applied CycleGAN [38] to generate fake satellite images, coupled with an additional SAFA module [39] to reduce discrepancies. However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise. Furthermore, the non-end-to-end training framework employed significantly increases training time and computational costs, reducing the efficiency and practical applicability. Additionally, existing work [5] utilized frozen foundation models (FMs) [40, 41] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different point of view. However, this method heavily relied on the ability of FMs for feature extraction. When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected. To reduce reliance on the feature extractor, CDIKTNet [42] trained with a small amount of paired UAV-satellite images to optimize the initial feature distribution, and then performed clustering to generate pseudo-labels, thereby improving the quality of pseudo-labels generation. However, this method inherently depends on the feature extractor, as it still requires limited paired UAV-satellite images for effective learning.
@@
II-C Self-Supervised Person Re-Identification
@@
ReID [43, 44] aims at matching the same person image captured by non-overlapping cameras. Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress. However, these achievements are facilitated by extensive human-labeled data. To address this limitation, self-supervised learning methods [6, 7] have been proposed. Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training. In addition, benefiting from the introduction of instance-level contrastive learning [45] with memory banks and momentum update strategies [46], many methods [47, 48] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations. However, these instance-level features was modified largely during training, resulting in unstable representations [49]. To address this issue, recent studies [50, 51] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations. Cluster Contrast [17] addresses the problem of cluster inconsistency by performing clustering and introducing a distinct representation for each cluster. Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features. ADCA[18] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment. Building on this, SDCL [52] employs a dual-path Transformer for shallow-to-deep contrastive learning, and refines pseudo-labels via consistency constraints across feature hierarchies, leading to more stable and discriminative cross-modal representations.
The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for UVGL. However, UVGL suffers from more severe viewpoint and scale variations. Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, severely degrading clustering effectiveness. Thereby, we incorporate instance-level and cluster-level features into a self-supervised UVGL method to learn more discriminative and robust cross-view representations.
III METHOD
@@
As illustrated in Fig. 3, the proposed framework is composed of two key modules, DHML and ICEL. Additionally, a PLE strategy is employed to optimize the self-supervised learning process. Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with cluster memory and a dual-path contrastive learning strategy to learn intra-view feature representations. To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory, thereby enhancing the discriminability of intra-view feature representations. Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints. By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form unified and consistent representations.
@@
Problem Formulation: In the UVGL, we are given a set of paired UAV-satellite images, denoted by . Here, represents UAV-view images and represents satellite-view images. Unlike the supervised setting, our self-supervised method is developed without using paired UAV-satellite images, where the correspondences between UAV-view and satellite-view images are entirely unavailable. Thereby, we randomly sample UAV and satellite images independently during training. The objective is to discover reliable cross-view correspondences under such conditions and accomplish the UVGL task.

### v3

# Without Paired Labeled Data: End-to-End Self-Supervised Learning for Drone-view Geo-Localization

@@ Abstract
Drone-view Geo-Localization (DVGL) aims to achieve accurate localization of drones by retrieving the most relevant GPS-tagged satellite images. However, most existing methods heavily rely on strictly pre-paired drone-satellite images for supervised learning. When the target region shifts, new paired samples are typically required to adapt to the distribution changes. The high cost of annotation and the limited transferability of these methods significantly hinder the practical deployment of DVGL in open-world scenarios. To address these limitations, we propose an end-to-end self-supervised learning method with a shallow backbone network. It employs a clustering algorithm to generate pseudo-labels and adopts a dual-path contrastive learning framework to learn discriminative intra-view representations. Furthermore, our method incorporates two core modules, including the dynamic hierarchical memory learning module (DHML) and the information consistency evolution learning module (ICEL). The DHML combines short-term and long-term memory to enhance intra-view feature consistency and discriminability. Meanwhile, the ICEL module utilizes a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, consequently improving cross-view feature alignment. To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced to enhance the quality of pseudo supervision. Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods. Our code is available at https://github.com/ISChenawei/DMNIL.
@@
Drone-view Geo-Localization, Self-Supervised Learning, Contrastive Learning.I Introduction
@@
Drone-View Geo-Localization (DVGL) is intended to achieve high-precision localization of drones in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to drone aerial images [1]. This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response. In recent years, existing DVGL methods have mainly relied on supervised learning using strictly paired drone-satellite images [2, 3]. Despite the promising performance demonstrated by these supervised methods, several intrinsic limitations are exposed. Firstly, the acquisition of instance-level paired drone-satellite images incurs additional training expenses [4]. Secondly, large amounts of unpaired cross-view image resources remain under-exploited. Lastly, when these methods are applied to new scenarios, the need arises for re-collection and re-pairing of drone-satellite images for training purposes to accommodate domain shifts. These limitations significantly hinder the practical deployment of DVGL in open-world scenarios. Thereby, it is necessary to explore a self-supervised method that can extract latent cross-view correlations from large-scale unpaired drone-satellite images.
In recent years, self-supervised person re-identification (ReID) methods [5, 6] have achieved remarkable progress to offer crucial theoretical foundations for DVGL. Typically, these methods employ clustering algorithms [7, 8] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning. However, as illustrated in Fig. 1, compared to conventional ReID tasks, DVGL presents more severe challenges, such as large intra-view ambiguities and cross-view discrepancies. As shown in Fig. 2, these issues hinder the backbone network from extracting stable and discriminative features at the early stage, which directly undermines the quality of pseudo-labels generated during the initial clustering phase. Without effective strategies to guide the learning of intra-view and cross-view consistency as well as feature discrimination, the model lacks the necessary ability to extract meaningful representations from low-quality pseudo-labels. As a result, the learning process may precipitate a detrimental feedback loop, wherein inaccurate pseudo-labels undermine the quality of feature representations to further compromise pseudo-label quality in subsequent iterations, ultimately leading to the collapse of self-supervised training. Although a current self-supervised method [9] proposed a coarse-to-fine framework for DVGL with a similar dual-path contrastive learning framework, it relies on explicit appearance-level transformations rather than intrinsically modeling the underlying cross-view feature relationships. In the shared space, it still exhibits limited cross-view alignment and susceptibility to feature noise.
@@
In this paper, we propose a novel end-to-end self-supervised learning method for DVGL. A lightweight and shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels. Combined with a dual-path contrastive learning strategy [17, 18], the backbone network coupled with a downstream clustering process forms the baseline network of our method for learning intra-view feature representations. Building upon this baseline, we design the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method. The DHML module captures the relationship between short-term and long-term memory to enhance the discriminability of intra-view feature representations. The ICEL module constructs a neighborhood-driven dynamic constraint mechanism by integrating neighborhood consistency constraints with mutual information optimization to facilitate cross-view feature alignment and discriminative representation learning. DHML enhances intra-view discrimination and ICEL improves cross-view consistency. Through joint optimization of intra-view and cross-view features, they collectively strengthen the feature discriminability and consistency. Additionally, a pseudo-label enhancement (PLE) strategy is employed to improve the reliability of pseudo-labels and guide more stable self-supervised feature learning. The main contributions can be summarized as follows.
@@
We propose an end-to-end self-supervised learning method for DVGL. Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to enhance feature consistency and discriminability for both intra-view and cross-view representation learning. Additionally, a pseudo-label enhancement (PLE) strategy is employed to guide more stable self-supervised learning.
@@
To address the intra-view ambiguity caused by spatial continuity across geographic viewpoints, we propose a dynamic memory hierarchical learning (DHML) module that captures the dependencies between different memory levels to adaptively learn the dynamics of intra-class and inter-class features within the intra-view.
@@
To address cross-view discrepancies and inconsistencies, we propose a novel information consistency evolution learning (ICEL) module which combines neighborhood consistency and mutual information optimization to learn consistent and discriminative cross-view representations, bridging the discriminability in intra-view learned by the DHML module to form a synergistic learning system.
@@
Extensive experiments demonstrate that our method, which employs a lightweight backbone, outperforms existing self-supervised methods on three DVGL benchmarks and even surpasses several state-of-the-art supervised methods.
@@
The remainder of this paper is organized as follows. Section II systematically reviews previous research. In Section III, the proposed self-supervised learning method is presented in detail. The experimental results are reported and analyzed in Section IV. Finally, conclusions are outlined in Section V.
@@ II Related Works
This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised learning method for DVGL. DVGL is typically regarded as a sub-task of cross-view geo-localization (CVGL). To provide a more comprehensive understanding of the field, we first review the overall progress in CVGL, including both supervised and self-supervised methods. Then, we summarize the main methods in self-supervised ReID that have provided analogous problem-solving paradigms.
@@ II-A Supervised Cross-View Geo-localization
CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images. Early research focused mainly on the CVGL in ground-view [19] with several benchmark datasets, such as CVUSA [20], CVACT [21], and VIGOR [22]. To address the challenge of ground-view CVGL, researchers have proposed various methods. CVM-Net [23] employed NetVLAD [24] to encode images as global descriptors to reduce visual discrepancies caused by viewpoint variations. TransGeo [25] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization. GeoDTR [26] leveraged a Transformer-based [27] feature extractor to disentangle geometric features, effectively addressing perspective shifts. Sample4Geo [28] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
With the continuous development and application of drone technology, DVGL has gradually become an important research direction and has achieved significant progress [29, 30, 31]. Several benchmark datasets for DVGL have been proposed, such as University-1652 [12], SUES-200 [32], and DenseUAV [33], which have driven further research in this field. In this scenario, researchers have proposed a variety of methods. LPN [34] extracted fine-grained features through local pattern partitioning to improve the ability to capture local details. IFS [35] utilized a multi-branch strategy to effectively fuse global and local features. CAMP [29] improved the accuracy of cross-view matching by contrastive attribute mining and position-aware partitioning, while DAC [3] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints. MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve the representation and alignment of features.
Nevertheless, the remarkable performances exhibited by these DVGL methods require extensive pre-processed drone-satellite image pairs. In this work, we concentrate on self-supervised learning approaches for DVGL, particularly addressing scenarios where labeled pairings are unavailable, which is crucial for real-world applications.
@@ II-B Self-Supervised Cross-View Geo-localization
Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for the real world, researchers have begun to explore self-supervised learning methods for CVGL. The work in [36] proposed a self-supervised learning method that used correspondence-free projection to transform ground panorama images into bird’s eye view images. It also applied CycleGAN [37] to generate fake satellite images, coupled with an additional SAFA module [38] to reduce discrepancies. However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise. Furthermore, the non-end-to-end training method employed significantly increases training time and computational costs with low practical applicability. Additionally, existing work [4] utilized frozen foundation models (FMs) [39] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different points of view. However, this method heavily relied on the ability of FMs for feature extraction. When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected. To reduce reliance on the feature extractor, CDIKTNet [40] trained with a small amount of paired drone-satellite images to optimize the initial feature distribution and then performed clustering to generate high-quality pseudo-labels. However, this method inherently depends on the feature extractor, as it still requires limited paired drone-satellite images for effective learning.
@@
II-C Methodological Parallels from Self-Supervised ReID
@@
ReID [41] aims to match the same person image captured by non-overlapping cameras. Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress. It has provided a fully end-to-end framework to eliminate the need for additional training processes and completely remove the reliance on feature extractor initialization, which offers strong inspiration for our consequent work in the DVGL task. However, these achievements are facilitated by extensive human-labeled data. To address this limitation, self-supervised learning methods [5, 6] have been proposed. Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training. In addition, benefitting from the introduction of instance-level contrastive learning [42] with memory banks and momentum update strategies [43], many methods [44, 45] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations. However, these instance-level features were largely modified during training, which resulted in unstable representations [46]. To address this issue, recent studies [47, 48] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations. Cluster-Contrast [17] contrasts instance features with these cluster-level representations to address the issue of inconsistent instance discrimination. Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features. ADCA [18] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment. Building on this, SDCL [49] employed a dual-path Transformer for shallow-to-deep contrastive learning, and refined pseudo-labels via consistency constraints across feature hierarchies to achieve more stable and discriminative cross-modal representations. The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for DVGL. However, DVGL suffers from more severe viewpoint and scale variations. Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, which would severely degrade clustering effectiveness.
A cluster-based self-supervised framework, which is developed concurrently with our work, is proposed with a similar ReID framework that employs intermediate state transition and local-to-global matching for cross-view alignment to mitigate viewpoint discrepancies. However, this method relies on explicit appearance-level transformations rather than intrinsically modeling cross-view semantic features, and lacks end-to-end optimization and uncertainty-aware pseudo-label refinement, making it sensitive to noise and limiting its adaptability. In this work, we integrate instance-level and cluster-level features in a self-supervised method for DVGL by employing reliable neighbor learning to facilitate implicit cross-view alignment and improve the discriminative power and robustness of the learned representations.
III End-to-End Self-Supervised Learning Method
@@
As illustrated in Fig. 3, the proposed method is composed of two key modules, DHML and ICEL. Additionally, a PLE strategy is employed to optimize the self-supervised learning process. Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with cluster memory and a dual-path contrastive learning strategy to learn intra-view feature representations. To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory. DHML module enables efficient discriminability of intra-view feature representations. Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints. By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form consistent and unified representations.
@@
Problem Formulation: In DVGL, a set of paired drone-satellite images is denoted as , where represents the drone-view image and represents the satellite-view image. Unlike the supervised setting, our self-supervised method is developed without using paired drone-satellite images, where the correspondences between drone-view and satellite-view images are entirely unavailable. The objective is to discover reliable cross-view correspondences under such conditions and to accomplish the DVGL task.